An aqueous solution started at pH= 4.58 but with the addition of an acid reduced to 1.96. At the end of the total V= 124.55 ml. Solid Li2CO3 was added (MM= 73.89G/mol) to reestablish the original pH.
What mass of bass is required
1
Expert's answer
2020-12-06T10:26:55-0500

Mass of "Li_2CO_3=x"

Molar mass of "Li_2CO_3=73.89g\/mol"

Volume of solution"=0.1245L"

"K_w=1.0\times 10^{-4}"

pH of solution"=1.96"

pH="-log [H^+][H^+]=10^{-pH}"

"10^{-4}=1.0\times 10^{-4}"

Dissociation equation is;

"Li_2CO_3\leftrightharpoons Li^{2+}+CO^{2-}_3"

"1.0\times 10^{-4}x=" "K_a" "=[Li^{2+}][CO_3^{2-}]\over Li_2CO_3"

"4.6\times 10^{-4}(1.96)\over 1.0\times 10^{-4}" "=9.016M"

"[CO_3^{2-}]=9.016M"

Mass of "Li_2CO_3=?"

"Mass=" "Mol\over L" "]CO_3^{2-}"

"=" "9.016mol\over L" "\times 0.1245L"

"=1.1225CO_3^{2-}"

"mLi_2CO_3=1.1225 mol\times 73.89g/mol=82.94g"

"\therefore" Mass of "Li_2CO_3=82.94g"